OpenStudy (rosales0228):

WILL GIVE MEDAL AND FAN!!! f(x) = (x + 2)2 and g(x) = x2 − (x + 2)2. Find f(x)-g(x)

OpenStudy (rosales0228):

2x2 + 8x + 8 3x2 + 8x + 8 x2 + 8x + 8 4x + 4 x2 − 8x + 8

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x) = (x + 2)^2}\) \(\normalsize\color{blue}{ g(x) = x^2 − (x + 2)^2}\) \(\normalsize\color{blue}{ f(x)~~-~~g(x)~ =(x + 2)^2~~-~( x^2 − (x + 2)^2~~)}\)

OpenStudy (solomonzelman):

(x+2)² - [ x² - (x+2)² ] (x+2)² - x² + (x+2)² 2(x+2)² - x² 2(x²+4x+4) - x² 2x²+8x+8 - x² x²+8x+8

OpenStudy (solomonzelman):

A parabola, x²+8x+8 x²+8x+16 - 8 (x+4)² - 8

OpenStudy (solomonzelman):

Well, you just need the `x²+8x+8`